Derby Phones is considering the introduction of a new model of headphones with the following price and cost characteristics: $ 270 per unit Sales price Variable costs Fixed costs 120 per unit 300,000 per month Assume that the projected number of units sold for the month is 5,000. Consider requirements (b), (c), and (d) independently of each other. Required a. What will the operating profit be? Operating proft b. What is the impact on operating profit if the sales price decreases by 10 percent? Increases by 20 percent? Sales price decreases by 10 percent: Operating profit by Sales price increases by 20 percent: Operating profit by

media%2F5a7%2F5a73eaa3-3e93-4c71-9643-a4
Answer #1

Solution:-

(a).Operating profit :-

Sales

= 5,000 * 270

= $1,350,000

Variable costs

= 5,000 * 120

= $600,000

Contribution margin = sales - variable cost

= 1,350,000 - 600,000

= $750,000

Fixed cost $300,000
Operating profit = contribution margin - fixed cost

= 750,000 - 300,000

= $450,000

(b).Operating profit when selling price decreased by 10% and increased by 20% :-

sales price by 10% = 270 - (270 * 10%)

= 270 - 27

= $243

sales price by 10% = $243

sales price by 20%= 270 + (270 *20%)

= 270 + 54

= $324

sales price by 20% = $324

selling price decreased by 10% selling price increased by 20%
Sales

= 5,000 * 243

= $1,215,000

= 5,000 * 324

= $1,620,000

Variable costs

= 5,000 * 120

= $600,000

= 5,000 * 120

= $600,000

Contribution margin =( sales - variable cost )

= 1,215,000 - 600,000

= $615,000

= 1,620,000 - 600,000

= $ 1,020,000

Fixed costs $300,000 $300,000
Net Operating profit = contribution marin - fixed cost

= 615,000 - 300,000

= $315,000

= 1,020,000 - 300,000

= $720,000

  • Here the selling price is decreased by 10% .then ,

operating profit = 450,000 - 315,000

= $135,000

Decreased operating profit = $135,000

  • Here selling price is increased by 20% .then

operating profit = 720,000 - 450,000

= $270,000

Increased operating profit = $270,000.

(C).Variable cost decreased by 10% and increased by 20%:-

variable cost decreased by 10% = 120 - (120 * 10%)

= 120 - 12

= $108

variable cost decreased by 10% = $108

variable cost increased by 20% = 120 + (120 * 20% )

= 120 + 24

= $144

variable cost increased by 20% = $144

selling price decreased by 10% selling price increased by 20%
Sales

= 5,000 * 270

= $1,350,000

= 5,000 * 270

= $1,350,000

Variable costs

= 5,000 * 108

= $540,000

= 5,000 * 144

= $720,000

Contribution margin =( sales - variable cost )

= 1,350,000 - 540,000

= $810,000

= 1,350,000 - 720,000

= $630,000

Fixed costs $300,000 $300,000
Net Operating profit = contribution marin - fixed cost

= 810,000 - 300,000

= $510,000

= 630,000 - 300,000

=$330,000

  • Here the variable cost is decreased by 10% .then ,

operating profit = 510,000 - 450,000

= $60,000

Decreased operating profit = $60,000

  • Here thevariable cost is increased by 20% .then ,

operating profit = 450,000 - 330,000

= $120,000

increased operating profit = $120,000

(d).

Fixed cost decreased by 20%:-

Fixed cost decreased by 20% = 300,000 - (300,000 * 20%)

= 300,000 - 60,000

= $240,000

Fixed cost decreased by 20% = $240,000

variable cost increases by 10% = 120 + (120 * 10%)

= 120 + 12

= $132

variable cost increases by 10% = $132

Sales

= 5,000 * 270

= $1,350,000

variable cost

= 5,000 *132

= $660,000

Contribution margin =( sales - variable cost )

= 1,350,000 - 660,000

= $690,000

Fixed cost $240,000
Net Operating profit = contribution marin - fixed cost

= 690,000 - 240,000

= $450,000